OpenStudy (anonymous):

Type A coffee costs Latoya $4.30 per pound, and type B coffee costs $5.50 per pound. This month, Latoya made 150 pounds of the blend, for a total cost of $708.60 . How many pounds of type A coffee did she use?

OpenStudy (igreen):

We can write a system of equations.. 4.30a + 5.50b = 708.6 a + b = 150 Where a = Type A coffee and b = Type B coffee

OpenStudy (igreen):

We can rearrange the 2nd equation and solve it for 'a', and solve for 'b'. a + b = 150 Subtract 'b' to both sides: a = -b + 150 Now we can plug in -b + 150 for 'a' in the 1st equation: 4.30a + 5.50b = 708.6 4.30(-b + 150) + 5.50b = 708.6 Distribute 4.30 into the parenthesis: -4.30b + 645 + 5.50b = 708.6 Can you solve the rest of that? @mbspires

OpenStudy (anonymous):

53?

OpenStudy (igreen):

No..actually we would add: -4.30b + 5.50b Can you add those?

OpenStudy (anonymous):

1.2b

OpenStudy (igreen):

Yes! That gives us: 1.2b + 645 = 708.6 Now we subtract 645 to both sides, what's 708.6 - 645?

OpenStudy (igreen):

Oh..you already solved all of it and got 53..

OpenStudy (igreen):

Nvm, you're correct.

OpenStudy (igreen):

Therefore b = 53 So now we plug that into any of the two equations. a + b = 150 Plug in 53 for 'b': a + 53 = 150 Now solve that for 'a'.

OpenStudy (anonymous):

97! thanks

OpenStudy (igreen):

Yep, that's your answer!